# Understanding Trends, Patterns, and Dynamics in Global Acquisitions: A Network Perspective

### v2

@@ 3.2.6 Closeness centrality
The closeness centrality calculates the mean distance from a node to other nodes in a network. The formula for closeness centrality is as follows:
@@
where n indicates the count of nodes accessible from node $u$, and $d(v,u)$ represents the shortest distance between nodes $v$ and $u$ [37]. We calculate this measure to determine which cities are closer to other cities in the cross-city acquisition network.
@@ 3.2.7 Weighted degree centrality
The weighted degree centrality calculates the sum of the weights of the edges connected to a node in a network. The formula for weighted degree centrality is as follows:
@@
where node $v$ is a neighbor of node u, and $w(v,u)$ is the weight of the edge between nodes $v$ and $u$ [37]. We calculate this metric to assess which cities acquire a greater number of organizations from other cities in the cross-city acquisition network.
@@ 3.2.8 PageRank centrality
The PageRank centrality calculates the ranking of a node within a network by considering the incoming link structure. The formula for PageRank centrality is as follows:
@@
where $A_{ij}$ is an element of the adjacency matrix, $d_{j}^{out}$ denotes the out-degree of node $j$, and $\alpha$ and $\beta$ are positive constants [37]. We calculate this metric to identify influential cities in the cross-city acquisition network in terms of their incoming link counts.
@@ 3.2.9 Eigenvector centrality
The eigenvector centrality calculates the centrality of a node considering the centrality of its neighboring nodes. It is determined by solving the equation
@@
where $A$ represents the adjacency matrix of the network and $x$ is the eigenvector. The eigenvalue associated with $A$ influences this calculation [37]. We employ this metric to assess the transitive influence among cities in the cross-city acquisition network.
@@ 3.2.10 Authority centrality
The authority centrality of a node within a network is derived from the principal eigenvector of the matrix $A^{T}A$, where $A$ stands for the network’s adjacency matrix [39]. We utilize this measure to assess the most influential countries in the cross-border acquisition network.
4 Results and discussion
@@ 4.1 Acquisition network
In this part, we delineate the construction process of the acquisition network using Crunchbase data. In this network, nodes represent organizations engaged in acquisitions. We establish directed edges from organization $A$ to organization $B$ if $A$ has acquired $B$, resulting in a directed network comprising 119,263 nodes and 92,788 edges. Considering the network’s extensive size, we selected a sample for visualization purposes. To accomplish this, we displayed the four largest Louvain communities from the largest weakly connected component of the network. Figure 7 illustrates this subgraph, encompassing 1,445 nodes and 1,447 edges, highlighting four distinct communities within the network.
Figure 7: A sampled subgraph of the acquisition network. Nodes are colored based on their community detected via the Louvain algorithm [40].
@@ 4.2 Common acquirer network
In this part, we delve into the common acquirer network, which is a projection of the acquisition network onto acquiree nodes. To construct this network, we establish an edge between two organizations whenever a common entity has acquired both of them. This results in an undirected weighted network, where the weight of an edge signifies the number of shared acquirers between two organizations. This network encompasses 68,297 nodes and 639,064 edges. Figure 8 depicts a sample of this network, showcasing the six largest Louvain communities within the largest connected component of the network. The resulting sample consists of 3,937 nodes and 44,340 edges. As illustrated in Figure 8, each community within the network is highlighted with a specific color. Companies within each community primarily originate from the United States. Table 1 presents the percentage of the major industry category within each community. As shown, most communities have Health Care or Manufacturing as their primary industry category, aligning with our previous observation indicating that these industries are the most common category groups among startups.
Figure 8: A sampled subgraph of the common acquirer network. Nodes are colored based on their community detected via the Louvain algorithm.
Table 1: Percentage of major industry categories within communities.
@@ 4.3 Common acquiree network
In this part, we describe the construction of the common acquiree network, which is another projection of the acquisition network. To build this network, we establish an undirected edge between two organizations if they have acquired the same organization. The weight of these edges corresponds to the number of shared acquirees between the two organizations. This process results in a weighted network comprising 4,063 nodes and 3,115 edges when applied to the entire dataset. To facilitate visualization, we have selected the eight largest communities from within the largest connected component, creating a sample network composed of 496 nodes and 593 edges. Figure 9 depicts this visualization with communities detected using the Louvain algorithm. Each community mainly comprises acquirer companies from the United States and the Financial Services industry. Table 2 represents the most common regions within each community, along with their respective percentages. According to this table, New York and England are the major regions from which acquirer companies originate, aligning with prior studies that highlight the prominence of these regions in startup ecosystems [32, 41].
Figure 9: A sampled subgraph of the common acquiree network. Nodes are colored based on their community detected via the Louvain algorithm.
Table 2: Percentage of major regions within communities.
@@ 4.4 Structural analysis
In this part, we present a summary of the structural properties of the acquisition network, common acquirer network, and common acquiree network. Based on the data in Table 3, it is evident that these networks have low densities, suggesting the likelihood of organizations forming smaller, distinct communities within these networks. The transitivity of the acquisition network is remarkably low, while that of the common acquirer network is close to 1. The common acquiree network exhibits a transitivity that is relatively low, but it is higher compared to that of the acquisition network. This outcome was anticipated since projection networks typically manifest strong community structures, resulting in high transitivity [42].
@@
We computed the average clustering coefficient and average shortest path length within the largest weakly connected component of the acquisition network. These metrics indicate that this network is not well-connected. The low average clustering coefficient suggests the presence of weak ties, and on average, it takes at least 14 steps to reach from one randomly selected organization to another. Conversely, the common acquirer network exhibits a contrasting scenario. It boasts a high average clustering coefficient and a small average shortest path length. This pattern aligns with the network’s configuration of well-connected clusters of organizations sharing common acquirers, as anticipated. On the other hand, the common acquiree network has a moderate average clustering coefficient and an average shortest path length close to that of the common acquirer network.
@@
For community detection, we employed the Infomap algorithm [43] to identify communities within these networks. The substantial number of communities implies acquisitions occurring among distinct subgroups of organizations. Additionally, we calculated the sizes of the largest connected components (both weakly and strongly connected for the acquisition network) in these networks. Notably, the presence of a strongly connected component with a size of three in the acquisition network indicates that at most three organizations are mutually involved in acquisitions. Conversely, the substantial size of the largest connected component (and the largest weakly connected component for the acquisition network) underscores the robust and centralized structures inherent in these networks.
Table 3: Structural properties of the acquisition network and its projections.
@@
Density $6.524\times 10^{-6}$ $2.740\times 10^{-4}$ $3.775\times 10^{-4}$
Transitivity $6.024\times 10^{-5}$ 0.993 0.081
Average clustering coefficient 0.002 0.828 0.064
Average shortest path length 14.344 7.242 6.825
Number of communities 32250 12786 1366
Largest weakly connected component size 25285 - -
Largest strongly connected component size 3 - -
Largest connected component size - 10427 1098
@@
Table 4: The most central nodes within the acquisition network and its projections.
@@ 4.5 Assortativity analysis
In this part, we calculate and interpret assortativity coefficients for the acquisition network, common acquirer network, and common acquiree network. Table 5 provides an overview of these coefficients for different node features. As shown, the acquisition network exhibits the highest assortativity coefficient with respect to the country feature. This suggests that organizations are more likely to pursue acquisitions within their own country, consistent with previous studies that highlight a higher inclination among companies to engage in domestic acquisitions [45]. Such acquisitions are known to yield greater returns for acquirers [46]. The next feature with the highest assortativity is the category group, indicating organizations’ preference for acquiring those in the same industries as themselves, aligning with prior studies [47, 26]. The third highest coefficient corresponds to the founding date of the organizations, suggesting that organizations involved in acquisitions tend to have close founding dates, consistent with previous research [26]. Notably, all coefficients are positive, indicating a prevailing inclination among organizations to acquire entities similar to themselves, aligning with findings from prior studies [47].
@@
Table 5: Assortativity coefficients of the acquisition network and its projections.
@@ 4.6 ERGM analysis
In this part, we present the ERGMs calculated for the acquisition network, common acquirer network, and common acquiree network. Our models utilize similar sets of node features and “edges” as parameters. Table 6 provides a summary of the ERGM statistics. According to the results, for these networks, all predictors are statistically significant $(p<0.0001)$ with positive estimates, except for the founding date, indicating that organizations with similar values for each feature are more likely to be connected. For instance, organizations within the same country are more likely to share the acquirer-acquiree relationship. This aligns with prior research indicating that companies tend to purchase firms located in closer geographic proximity [50]. Conversely, negative estimations regarding the founding date suggest that this feature cannot be the cause of the connection between organizations within the networks. Furthermore, both networks have positive estimates for “edges”, indicating that they have higher densities than random networks with similar numbers of nodes. In other words, nodes in both networks are more inclined to be connected to each other [51]. The negative estimate for “edges” indicates that the common acquiree network exhibits lower density in comparison to random networks possessing similar topological features. The AIC values for each model, as shown in Table 6, reveal that the ERGM of the common acquiree network has a better fit (lower AIC value) compared to other networks.
Table 6: Estimates from the ERGMs of the acquisition network and its projections.
@@
Founding date (year-month) $***-5.298\times 10^{-5}***$ $(5.048\times 10^{-7})$ $***-4.678\times 10^{-5}***$ $(2.647\times 10^{-7})$ $***-9.727\times 10^{-6}***$ $(2.819\times 10^{-6})$
@@ 4.7 Temporal analysis
In this part, we investigate the evolution of the acquisition network over time. We begin our analysis with the January 2000 network and subsequently consider one-month intervals between consecutive time points, resulting in 283 time points up to the present data collection date. At each time point, we compute various structural metrics for the corresponding acquisition network.
@@
Figure 10 illustrates the temporal changes in network density. The visual representation reveals a declining trend, suggesting that the network has gradually become sparser over time. In essence, organizations within the network have exhibited reduced connections over time. This trend could be attributed to the increase in the number of investors and venture capitalists, and the fact that we are no longer observing a dominant monopoly [52, 53].
Figure 10: Time series of the density of the acquisition network.
@@
Figure 11 illustrates the average clustering coefficient of the network at different time points. According to this figure, the network’s average clustering coefficient remained at zero before 2012, indicating a lack of connections between the organizations’ neighborhoods within the network. However, starting from 2012, it shows an upward trend, reaching its peak in 2013. Subsequently, there is a decline observed until 2014, followed by a general increasing trend with intermittent fluctuations. This pattern suggests a gradual densification of local connections over time, even as the overall connections have become sparser. It could be related to the European debt crisis in the early 2010s, which might have prompted many companies to attempt acquiring neighboring businesses as a strategic move to expand and overcome economic challenges [54].
Figure 11: Time series of the average clustering coefficient of the acquisition network.
@@
Figure 12 illustrates the average shortest path length of the acquisition network over time. The metric remains around 2 until 2007, after which it steadily increases, reaching its initial peak in the same year. Subsequently, from 2007 onward, it gradually ascends to its second peak in 2015. Following this, it displays a gradual decrease and eventually fluctuates around 14. Essentially, this suggests that in recent years, on average, organizations required approximately 14 hops to reach another organization within the network. This pattern could be related to the 2007–2008 financial crisis, which might have resulted in the fragmentation of the acquisition network in the subsequent years [55].
Figure 12: Time series of the average shortest path length of the acquisition network.
@@
Figure 13 displays the changes in the number of communities identified within the acquisition network. As illustrated, it follows an increasing pattern, indicating a rising level of segregation in the network. In other words, organizations within the network have formed more distinct groups, resulting in a more complex structure. This could be related to the increase in the number of industries in which acquisitions occur, considering that acquisitions mostly happen between companies within the same industry [56, 47, 26].
Figure 13: Time series of the number of communities identified within the acquisition network.
Figure 14 represents the number of weakly connected components in the network over time. Similar to the number of communities, the number of connected components has also increased. In fact, the network has become more fragmented and less cohesive, further highlighting the absence of a monopoly in acquisition transactions in recent years [52, 53].
Figure 14: Time series of the number of weakly connected components within the acquisition network.
@@
Figure 15 represents the evolution of a sample of the acquisition network over time. We can observe the aforementioned trends in these networks. For instance, the average clustering coefficient of the network has increased, while the density of the network has declined. Moreover, both the number of communities and the number of connected components have increased.
Figure 15: Evolution of a sample of the acquisition network over time.

### v3

@@ 3.4.6 Closeness centrality
We measure the closeness centrality to determine which cities are closer to other cities in the cross-city acquisition network. Closeness centrality is useful for our analysis as cities with high values for this measure can swiftly affect economic trends and business practices, and facilitate acquisition. This metric calculates the mean distance from a node to other nodes in a network. The formula for closeness centrality is as follows:
@@
where n indicates the count of nodes accessible from node $u$, and $d(v,u)$ represents the shortest distance between nodes $v$ and $u$ [50].
@@ 3.4.7 Weighted degree centrality
We measure the weighted degree centrality to assess which cities acquire a greater number of organizations from other cities in the cross-city acquisition network. Weighted degree centrality is particularly suitable for this research as it offers a more nuanced measure of a city’s influence by not only considering the number of its connections but also their strength. This metric calculates the sum of the weights of the edges connected to a node in a network. The formula for weighted degree centrality is as follows:
@@
where node $v$ is a neighbor of node u, and $w(v,u)$ is the weight of the edge between nodes $v$ and $u$ [50].
@@ 3.4.8 PageRank centrality
We measure the PageRank centrality to identify influential cities in the cross-city acquisition network in terms of their incoming link counts. PageRank centrality is especially fitting for our analysis since it not only considers the number of incoming links but also reflects the significance of the linking cities. This metric calculates the ranking of a node within a network by considering the incoming link structure. The formula for PageRank centrality is as follows:
@@
where $A_{ij}$ is an element of the adjacency matrix, $d_{j}^{out}$ denotes the out-degree of node $j$, and $\alpha$ and $\beta$ are positive constants [50].
@@ 3.4.9 Eigenvector centrality
We measure the eigenvector centrality to assess the transitive influence among cities in the cross-city acquisition network. Eigenvector centrality is suitable for this study as it considers not only the number of links but also connections to other highly influential cities. This metric calculates the centrality of a node considering the centrality of its neighboring nodes. It is determined by solving the equation
@@
where $A$ represents the adjacency matrix of the network and $x$ is the eigenvector. The eigenvalue associated with $A$ influences this calculation [50].
@@ 3.4.10 Authority centrality
We measure the authority centrality to assess the most influential countries in the cross-border acquisition network. This metric is specifically relevant to our analysis as it highlights countries that are not only well-connected but also recognized as authorities in acquisitions. The authority centrality of a node within a network is derived from the principal eigenvector of the matrix $A^{T}A$, where $A$ stands for the network’s adjacency matrix [42].
4 Results
@@ 4.1 Acquisition network
The acquisition network comprises 119,263 nodes and 92,788 edges. Considering the network’s extensive size, we selected a sample for visualization purposes. To accomplish this, we displayed the four largest Louvain communities from the largest weakly connected component of the network. Figure 8 illustrates this subgraph, encompassing 1,445 nodes and 1,447 edges, highlighting four distinct communities within the network.
Figure 8: A sampled subgraph of the acquisition network. Nodes are colored based on their community detected via the Louvain algorithm [7]. The percentage of startups in each community is specified next to the community ID.
@@ 4.2 Common acquirer network
The common acquirer network encompasses 68,297 nodes and 639,064 edges. Figure 9 depicts a sample of this network, showcasing the six largest Louvain communities within the largest connected component of the network. The resulting sample consists of 3,937 nodes and 44,340 edges. As illustrated in Figure 9, each community within the network is highlighted with a specific color. Companies within each community primarily originate from the United States. Table 1 presents the percentage of the major industry category within each community. As shown, most communities have Health Care or Manufacturing as their primary industry category, aligning with our previous observation indicating that these industries are the most common category groups among startups.
Figure 9: A sampled subgraph of the common acquirer network. Nodes are colored based on their community detected via the Louvain algorithm. The percentage of startups in each community is specified next to the community ID.
Table 1: Percentage of major industry categories within communities of the sample common acquirer network in Figure 9. The table highlights that Manufacturing and Health Care are the dominant industry categories in these communities.
@@ 4.3 Common acquiree network
The common acquiree network comprises 4,063 nodes and 3,115 edges. To facilitate visualization, we have selected the eight largest communities from within the largest connected component, creating a sample network composed of 496 nodes and 593 edges. Figure 10 depicts this visualization with communities detected using the Louvain algorithm. Each community mainly comprises acquirer companies from the United States and the Financial Services industry. Table 2 represents the most common regions within each community, along with their respective percentages. According to this table, New York and England are the major regions from which acquirer companies originate, aligning with prior studies that highlight the prominence of these regions in startup ecosystems [49, 34].
Figure 10: A sampled subgraph of the common acquiree network. Nodes are colored based on their community detected via the Louvain algorithm. The percentage of companies in each community is specified next to the community ID.
Table 2: Percentage of major regions within communities of the sample common acquiree network in Figure 10. The table highlights that New York and England are the prominent regions in these communities.
@@ 4.4 Structural analysis
Table 5, 5, and 5 represent a summary of the structural properties of the acquisition network, the common acquirer network, and the common acquiree network, respectively. We considered the Erdős–Rényi (ER) model [8] to create baseline models of each of these networks, which are useful for validating the observed topological features of the networks. These baseline models have the same number of nodes and probability of edge creation as their respective real networks. Based on the data provided in the tables, it is evident that these networks have low densities, suggesting the likelihood of organizations forming smaller, distinct communities within these networks. The transitivity of the acquisition network is remarkably low, while that of the common acquirer network is close to 1. The common acquiree network exhibits a transitivity that is relatively low, but it is higher compared to that of the acquisition network. This outcome was anticipated since projection networks typically manifest strong community structures, resulting in high transitivity [52].
@@
We computed the average clustering coefficient and average shortest path length within the largest weakly connected component of the acquisition network. These metrics indicate that this network is not well-connected. The low average clustering coefficient suggests the presence of weak ties, and on average, it takes at least 14 steps to reach from one randomly selected organization to another. Conversely, the common acquirer network exhibits a contrasting scenario. It boasts a high average clustering coefficient and a small average shortest path length. This pattern aligns with the network’s configuration of well-connected clusters of organizations sharing common acquirers, as anticipated. On the other hand, the common acquiree network has a moderate average clustering coefficient and an average shortest path length close to that of the common acquirer network. The transitivity and average clustering coefficient values of the random models are zero or close to zero, which is expected, as ER models cannot capture these common features of real networks [11]. The average shortest path lengths of the ER models for the acquisition network and the common acquiree network are higher than those of the networks themselves, indicating the presence of economic hubs that shorten the distance between companies. However, for the common acquirer network, the ER model exhibits a lower average shortest path length, which is in line with the fragmentation of the network into distinct clusters with fewer connections between them.
@@
Additionally, we calculated the sizes of the largest connected components (both weakly and strongly connected for the acquisition network) in these networks. Notably, the presence of a strongly connected component with a size of three in the acquisition network indicates that at most three organizations are mutually involved in acquisitions. Conversely, the substantial size of the largest connected component (and the largest weakly connected component for the acquisition network) underscores the robust and centralized structures inherent in these networks. The size of the largest strongly connected component of the ER model for the acquisition network is 5, which is higher than that of the real network. Moreover, in terms of the largest weakly connected component size, all the three networks exhibits lower values compared to their ER counterparts. These observations are expected as random models do not have community structure and their degree distributions are more uniform, resulting in the formation of large connected components.
Table 3: Comparison of structural properties between the acquisition network and its baseline model. This table reveals that the acquisition network has a decentralized structure comprising smaller, distinct communities.
@@
Acquisition network Baseline model
Density $6.524\times 10^{-6}$ $6.530\times 10^{-6}$
Transitivity $6.024\times 10^{-5}$ $2.738\times 10^{-5}$
Average clustering coefficient 0.002 0
Average shortest path length 14.194 23.1727
Largest weakly connected component size 25285 73669
Largest strongly connected component size 3 5
Table 4: Comparison of structural properties of the common acquirer network and its baseline model. This table illustrates that the common acquirer network has tighter interconnections.
Property Measurement
Common acquirer network Baseline model
Density $2.740\times 10^{-4}$ $2.740\times 10^{-4}$
Transitivity 0.993 $2.707\times 10^{-4}$
Average clustering coefficient 0.959 $2.695\times 10^{-4}$
Average shortest path length 7.217 4.0775
Largest connected component size 10427 68297
Table 5: Comparison of structural properties of the common acquiree network and its baseline model. This table indicates that the common acquiree network exhibits more closely connected nodes.
Property Measurement
Common acquiree network Baseline model
Density $3.775\times 10^{-4}$ $3.890\times 10^{-4}$
Transitivity 0.081 $5.896\times 10^{-4}$
Average clustering coefficient 0.108 $3.216\times 10^{-4}$
Average shortest path length 6.922 15.451
Largest connected component size 1098 2591
@@
Table 6: The most central nodes within the acquisition network and its projections. This table demonstrates that pivotal acquiring firms are from the United States, while European startups are central in acquisition transactions.
@@ 4.5 Assortativity analysis
Table 7 provides an overview of assortativity coefficients corresponding to various node features in the acquisition network, common acquirer network, and common acquiree network. As shown, the acquisition network exhibits the highest assortativity coefficient with respect to the country feature. This suggests that organizations are more likely to pursue acquisitions within their own country, consistent with previous studies that highlight a higher inclination among companies to engage in domestic acquisitions [13]. Such acquisitions are known to yield greater returns for acquirers [64]. The next feature with the highest assortativity is the category group, indicating organizations’ preference for acquiring those in the same industries as themselves, aligning with prior studies [60, 44]. The third highest coefficient corresponds to the founding date of the organizations, suggesting that organizations involved in acquisitions tend to have close founding dates, consistent with previous research [44]. Notably, all coefficients are positive, indicating a prevailing inclination among organizations to acquire entities similar to themselves, aligning with findings from prior studies [60].
@@
Table 7: Assortativity coefficients of the acquisition network and its projections. The data in this table suggest that companies from the same country or category, or with similar founding dates, tend to engage in acquisition transactions with each other.
@@ 4.6 ERGM analysis
Our ERGM models for the acquisition network, common acquirer network, and common acquiree network utilize similar sets of node features and “edges” as parameters. Table 8 provides a summary of the ERGM statistics. According to the results, for these networks, all predictors are statistically significant $(p<0.0001)$ with positive estimates, except for the founding date, indicating that organizations with similar values for each feature are more likely to be connected. For instance, organizations within the same country are more likely to share the acquirer-acquiree relationship. This aligns with prior research indicating that companies tend to purchase firms located in closer geographic proximity [22]. Conversely, negative estimations regarding the founding date suggest that this feature cannot be the cause of the connection between organizations within the networks. Furthermore, the acquisition network and the common acquirer network have positive estimates for “edges”, indicating that they have higher densities than random networks with similar numbers of nodes. In other words, nodes in both networks are more inclined to be connected to each other [48]. The negative estimate for “edges” indicates that the common acquiree network exhibits lower density in comparison to random networks possessing similar topological features. The Akaike Information Criterion (AIC) values for each model, as shown in Table 8, reveal that the ERGM of the common acquiree network has a better fit (lower AIC value) compared to other networks.
Table 8: Estimates from the ERGMs of the acquisition network and its projections. The data in this table show that companies with the same set of features, except for the founding date, are more likely to be connected to each other in these networks.
@@
Founding date (year-month) $-5.298\times 10^{-5}***$ $(5.048\times 10^{-7})$ $-4.678\times 10^{-5}***$ $(2.647\times 10^{-7})$ $-9.727\times 10^{-6}***$ $(2.819\times 10^{-6})$
@@ 4.7 Temporal analysis
To investigate the evolution of the acquisition network over time, we begin our analysis with the January 2000 network and subsequently consider one-month intervals between consecutive time points, resulting in 283 time points up to the present data collection date. At each time point, we compute various structural metrics for the corresponding acquisition network.
@@
Figure 11 illustrates the temporal changes in the density, average clustering coefficient, average shortest path length, and the number of weakly connected components of the network. The visual representation reveals a declining trend for density, suggesting that the network has gradually become sparser over time. In essence, organizations within the network have exhibited reduced connections over time. This trend could be attributed to the increase in the number of investors and venture capitalists, and the fact that we are no longer observing a dominant monopoly [28, 45].
@@
According to Figure 11, the network’s average clustering coefficient remained at zero before 2012, indicating a lack of connections between the organizations’ neighborhoods within the network. However, starting from 2012, it shows an upward trend, reaching its peak in 2013. Subsequently, there is a decline observed until 2014, followed by a general increasing trend with intermittent fluctuations. This pattern suggests a gradual densification of local connections over time, even as the overall connections have become sparser. It could be related to the European debt crisis in the early 2010s, which might have prompted many companies to attempt acquiring neighboring businesses as a strategic move to expand and overcome economic challenges [51].
@@
Based on this figure, the average shortest path length of the acquisition network remains around 2 until 2007, after which it steadily increases, reaching its initial peak in the same year. Subsequently, from 2007 onward, it gradually ascends to its second peak in 2015. Following this, it displays a gradual decrease and eventually fluctuates around 14. Essentially, this suggests that in recent years, on average, organizations required approximately 14 hops to reach another organization within the network. This pattern could be related to the 2007–2008 financial crisis, which might have resulted in the fragmentation of the acquisition network in the subsequent years [56].
@@
As illustrated in Figure 11, the number of weakly connected components follows an increasing pattern, indicating a rising level of segregation in the network. In fact, the network has become more fragmented and less cohesive, further highlighting the absence of a monopoly in acquisition transactions in recent years [28, 45]. In other words, organizations within the network have formed more distinct groups, resulting in a more complex structure. This could be related to the increase in the number of industries in which acquisitions occur, considering that acquisitions mostly happen between companies within the same industry [30, 60, 44].
Figure 11: Time series of density, average clustering coefficient, average shortest path length, and the number of weakly connected components within the acquisition network starting from 2000. These plots together suggest that the network has become more fragmented while comprising denser clusters.
@@
Figure 12 represents the evolution of a sample of the acquisition network over time. We can observe the aforementioned trends in these networks. For instance, the average clustering coefficient of the network has increased, while the density of the network has declined. Moreover, the number of connected components have increased.
Figure 12: Evolution of a sample of the acquisition network over time. This figure further confirms that the network has become sparser and more complex in terms of community structure.